OpenStudy (camerondoherty):

The function \[f(x) = -x^2 - 7x + 30\] shows the relationship between the vertical distance of a diver from a pool's surface f(x), in feet, and the horizontal distance x, in feet, of a diver from the diving board. What is a zero of f(x), and what does it represent? x = 10; the diver hits the water 10 feet away horizontally from the board. x = 3; the diver hits the water 3 feet away horizontally from the board. x = 10; the diver jumps in the pool at 10 feet per second. x = 3; the diver jumps in the pool at 3 feet per second.

OpenStudy (camerondoherty):

I just found the zeroes, I think it was asking for those in the question -10, 3 If I did my math right cx

OpenStudy (michele_laino):

please solve your quadratic equation, using the subsequent formula for finding the zeroes: \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] if the equation is: \[ax ^{2}+bx+c=0\]

OpenStudy (acxbox22):

you are correct ;)

OpenStudy (camerondoherty):

Do you want me to show my work, orrr...

OpenStudy (acxbox22):

solving for zeroes, 0=-x^2-7x+30 0=-(x+10)(x-3) x=-10,3 using the zero product theorum

OpenStudy (michele_laino):

I think that only x=3 is acceptable, since the situation described in your problem is: 

OpenStudy (acxbox22):

x cant be -10 in the problem because negative distance isnt possible

OpenStudy (camerondoherty):

Yes, that's what I thought too. Due to the scenario it couldn't be negative, and none of the options contain -10 either. cx

OpenStudy (camerondoherty):

So if x = 3, and x is the horizontal distance, in feet, of the distance of the diver from the diving board, then the horizontal distance is 3?

OpenStudy (acxbox22):

ya

OpenStudy (camerondoherty):

Oh ok, thanks both of you c: I wish I could give you both a medal, both of your responses were the best c:

OpenStudy (michele_laino):

thank you! @camerondoherty

OpenStudy (camerondoherty):

No, Thank You! c:

OpenStudy (acxbox22):

no problem :)

OpenStudy (michele_laino):

I give medal @acxbox22

OpenStudy (acxbox22):

thanks

OpenStudy (michele_laino):

you deserve it!

OpenStudy (ttp):

b? @acxbox22

OpenStudy (acxbox22):

? 1 month later??

OpenStudy (ttp):

yeah i'm just now getting to this problem :/

OpenStudy (acxbox22):

oh ya it is b from the work above

OpenStudy (ttp):

Okay got it!!! @acxbox22

OpenStudy (acxbox22):

yep ;)

OpenStudy (ttp):

thanks! @acxbox22 :)

OpenStudy (acxbox22):

it is a 2 in 1 problem :)

OpenStudy (ttp):

hahah yeah xD